How do I find the value for x? if 8 and 6 are the legs and x is the base.

Answer by stanbon(75887)

How do I find the value for x? if 8 and 6 are the legs and x is the base.
----
If x is the hypotenuse of a right triangle,
x^2 = 8^2+6^2
---
x^2 = 64+36
---
x^2 = 100
x = 10
-------------------------
Cheers,
Stan H.